Circular No. 7778

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001ij, 2001ik, 2001il, 2001im, 2001in G. Miknaitis, University of Washington, on behalf of the Sloan Digital

Sky Survey (SDSS), reports the detection of five apparent supernovae in images taken with the SDSS 2.5-m telescope on Nov. 21.3 UT. The supernovae were detected by subtraction against previous images taken by the SDSS 2.5m on 1998 Sept. 19.4 UT (available publicly as part of the SDSS Early Data Release). SNe 2001ij, 2001ik, and 2001il were confirmed in images taken by the SDSS on Nov. 11.3. SNe 2001il, 2001im, and 2001in were confirmed by R-band images taken by K. Krisciunas at the Cerro Tololo 1.5-m reflector on Dec. 12.5. SDSS g′ magnitudes are given in the table below for each supernova, as are the offsets from the center of each host galaxy; the respective r′ magnitudes for Nov. 21.3 were 19.8, 20.9, 20.6, 21.4, and 21.2. SN 2001ij is located in MCG +00-7-84.

SN  2001 UT  α_{2000}  δ_{2000}  g′  Offset
2001ij Nov. 21.3 2^{h}42^{m}31.^{s}63 − 1^{o}00′51′′.1 20.7 1′′.5 W, 4′′.9 N
2001ik Nov. 21.3 1 33 41.01 − 0 26 42.4 21.9 3′′.0 W, 0′′.4 N
2001il Nov. 21.3 2 26 41.35 − 0 57 31.7 20.4 1′′.5 E, 0′′.5 N
2001im Nov. 21.3 2 04 16.80 + 0 39 11.0 21.0 1′′.5 W, 1′′.3 N
2001in Nov. 21.3 1 01 51.11 + 0 41 21.7 22.7 7′′.5 W, 2′′.2 S

COMET P/2001 X3 (LINEAR) R. Huber, Lincoln Laboratory, Massachusetts Institute of Technology,

reports that an object discovered by LINEAR on Dec. 7 was found to be diffuse on Dec. 17 (with prediscovery LINEAR data back to Sept. 10). G. Hug, Eskridge, KS, reports that the object is probably diffuse on CCD images taken on Dec. 19.1 UT (red mag 17.2–17.8).

2001 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 7.08304  23^{h}18^{m}08.^{s}17  +5^{o}59′14′′.1  19.9

Full astrometry and orbital elements appear on MPEC 2001-Y15:

Epoch = 2002 Jan. 6.0 TT

T = 2001 Dec. 30.7555 TT  ω = 163.^{o}6343
e = 0.539153  Ω = 240.7107 }2000.0
q = 1.584056 AU  i = 13.4607

a = 3.437274 AU n^{o} = 0.1546617 P = 6.373 years

2001 December 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green